Dear Sir/Madam:
A pleasant day!I have a math problem here involving arithmetic sequence Given a sequence a1 , a2 , a3 ................. an and if a4=4 and a6=6 and an=an-1 + an-2 for n=1, 2, 3…… Find an

Answer by richwmiller(17219)

Do you mean An=A(n-1) + A(n-2)?

A4=A3+A2=4
A6=A5+A4=6
A5+4=6
A5=2
so we have
A4=4
A5=2
A6=6
A7=8
A8=14
A9=22
A10=36
A11=48
A12=84
That is not an arithmetic sequence.